# Universality of Estimator for High-Dimensional Linear Models with Block Dependency

### v1

@@ C.3. Proof of Theorem 12
All the conditions for the Theorem 4.1 in Thrampoulidis et al. (2018) trivially hold from Theorem 11 and assumptions of Theorem 12 as in Han and Shen (2023). ∎
Appendix D Proof of Auxiliary Results
D.1. Proof of Lemma 13
@@
Define $\bar{H}_{\beta}:=\beta\cdot\bar{H}$ and $F_{\beta}(A):=F_{\beta}(\bar{H}(w,A)_{w\in\mathcal{W}})$. As an initial step, we apply the result from Lemma 5, reducing the difference to the third derivatives of the functions involved. As shown in Appendix B.1, the third derivative of the function $g(n^{-1}F_{\beta}(\cdot))$ can be bounded by the derivatives of the loss function $\bar{H}$. For any $i\in\{1,\cdots,n\}$ and $j,k,\ell\in\{1,\cdots,d\}$, we have the following upper bound of derivatives of $\bar{H}$.
$|\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{2}_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{2}}\rangle_{\bar{H}_{\beta}})$
$|\langle\partial_{i,j}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{3}_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{3}}\rangle_{\bar{H}_{\beta}})$
where we define $\bar{H}^{(i)}(w,A):=\ell_{i:\rho}((Aw)_{i})\geq 0$. so $\bar{H}(w,A)=\sum_{i=1}^{n}\bar{H}^{(i)}(w,A)+nR(w)$. Therefore, any $i\in\{1,\cdots,n\}$ and $j,k,\ell\in\{1,\cdots,d\}$, we have
@@
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}})$
@@
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}})$
@@
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}})$
@@
$\lesssim L_{n}^{3}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}+q_{2}}\rangle_{\bar{H}_{\beta}})$
@@
$\lesssim L_{n}^{3}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}+q_{2}}\rangle_{\bar{H}_{\beta}})$
@@
$\leq L_{n}^{3}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{3}}\rangle_{\bar{H}_{\beta}}$
@@
$|\partial_{i,j}\partial_{i,k}\partial_{i,\ell}g(n^{-1}F_{\beta}(A))|\lesssim K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+\langle\bar{H}^{(i)}(w,A)^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}\right]$
where we define $\mathbf{q}=\max\{3q_{1},2q_{1}+2q_{2},q_{3}\}$. So, we obtain
$E[|\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}\partial_{i,(B_{j})_{\ell^{\prime\prime}}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))||X^{(B_{j})}_{i}]$
$\lesssim$ $K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}|X^{(B_{j})}_{i}]\right]$
By using Chebysev’s association inequality, we have
$E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}|X^{(B_{j})}_{i}]\leq E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]$
@@
$\begin{split}&E[|\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}\partial_{i,(B_{j})_{\ell^{\prime\prime}}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))||X^{(B_{j})}_{i}]\\ \lesssim&K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+E[\langle(\ell_{i;\rho}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]\right]\end{split}$ (19)
@@
$\begin{split}&E[\langle(\ell_{i;\rho}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]\\ &\leq E[\langle(\|\ell_{i;\rho}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i})-\ell_{i}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i})\|_{\infty}\\ &\quad+\ell_{i}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}]\\ &\lesssim_{q}(1+E[\langle(\ell_{i}(((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}])\\ &\lesssim_{q}(1+E[\langle L^{\mathbf{q}}_{\ell_{i}}(1+|((\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i})^{\top}w)_{i}|^{\mathbf{q}_{0}})\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}])\\ &\lesssim_{q}L^{\mathbf{q}}_{\ell_{i}}\left(1+E\left[\langle|((\mathbf{Z}_{i,j}^{0})^{\top}w)_{i}|^{\mathbf{q}_{0}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}\right]+E\left[\langle|t(X^{(B_{j})}_{i})^{\top}w|^{\mathbf{q}_{0}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}\right]\right)\end{split}$ (20)
where we define $\mathbf{q}_{0}=q_{0}\mathbf{q}$ and $\bar{\mathbf{q}}_{0}:=2\lceil\mathbf{q}_{0}/2\rceil$, respectively. By trivial calculation, we have
$|t(X^{(B_{j})}_{i})^{\top}w|^{\mathbf{q}_{0}}\leq t^{\mathbf{q}_{0}}L^{\mathbf{q}_{0}}_{n}\left(\sum_{m=1}^{|B_{j}|}|(X_{i}^{(B_{j})})_{m}|\right)^{\mathbf{q}_{0}}.$ (21)
@@
$\begin{split}E\left[\langle|((\mathbf{Z}_{i,j}^{0})^{\top}w)_{i}|^{\bar{\mathbf{q}}_{0}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}\right]\leq&E\left[\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\mathbf{q}_{0}}}\langle w_{k_{1}}\cdots w_{k_{\bar{\mathbf{q}}_{0}}}\rangle_{\bar{H}_{\beta;-i}}\right]\\ \leq&\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]E\left[\langle w_{k_{1}}\cdots w_{k_{\bar{\mathbf{q}}_{0}}}\rangle_{\bar{H}_{\beta;-i}}\right]\\ \leq&L_{n}^{\bar{\mathbf{q}}_{0}}\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\end{split}$ (22)
@@
$\begin{split}&E[|\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}\partial_{i,(B_{j})_{\ell^{\prime\prime}}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))|X^{(B_{j})}_{i}]\\ \lesssim&K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+L^{\mathbf{q}}_{\ell_{i}}L_{n}^{\bar{\mathbf{q}}_{0}}\left(1+\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]+t^{\bar{\mathbf{q}}_{0}}\left(\sum_{m=1}^{|B_{j}|}|(X^{(B_{j})}_{i})_{m}|\right)^{\bar{\mathbf{q}}_{0}}\right)\right]\end{split}$ (23)
holds. The second term expresses the sum of all combinations of $\mathbf{{\bar{q}}}_{0}$ random variables exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$. Combining Lemma 5 with (23), we attain
$\left|E\left[\int^{1}_{0}(1-t)^{2}\sum_{\ell=1}^{|B_{j}|}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))dt\right]\right|$
[…]
$\lesssim$ $K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$
$+$ $K_{g}n^{-1}\beta^{2}L_{n}^{\bar{\mathbf{q}}_{0}+3}D^{3}_{\ell}(\rho)L^{\mathbf{q}}_{\ell_{i}}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\left(1+\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}}+\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right)\right].$
@@ D.2. Proof of Lemma 14
We divide this summation into three parts based on $\ell$, i.e. (i)$11\leq\ell\leq\bar{\mathbf{q}}_{0}-$ 1, (ii) $\ell=\bar{\mathbf{q}}_{0}-\alpha$ and (iii) $(\bar{\mathbf{q}}_{0}+1)-\alpha\leq\ell\leq\bar{\mathbf{q}}_{0}$, and denote corresponding summation as $A_{1}$, $A_{2}$ and $A_{3}$, respectively.
@@ 3. Step 3
    4. Step 4
      Set $\gamma_{2,\ell}$ as the number of variables included in the same dependent index cell as $C^{j}_{ik_{(\gamma_{1,\ell})+1}}$. As in Step 2, the possible ways to select variables is $d^{(\gamma_{2,\ell})-1}$. Denote those variables as $C^{j}_{i,k_{2}},\cdots,C^{j}_{i,k_{\gamma_{1,\ell}}}$.
    5. Step 5
@@
    $N_{1}$ $\lesssim_{q}\sum_{\ell=1}^{\bar{\mathbf{q}}_{0}-\alpha-1}p^{\ell}d^{\sum_{\ell^{*}=1}^{\ell}(\gamma_{\ell^{*},\ell}-1)}$
    $=\sum_{\ell=1}^{\bar{\mathbf{q}}_{0}-\alpha-1}p^{\ell}d^{\bar{\mathbf{q}}_{0}-\ell}$
    where the last equality follows from the definition of $\gamma_{\ell^{*},\ell}$, i.e. $\sum_{\ell^{*}=1}^{\ell}\gamma_{\ell^{*},\ell}=\bar{\mathbf{q}}_{0}$. Moreover, because $d\leq p$ holds, for any $21\leq\ell\leq\bar{\mathbf{q}}_{0}-\alpha-$ 2, we have
@@
    $=$ $\sum_{\begin{subarray}{c}\ell_{1}=1\\ \ell_{1}\neq j\end{subarray}}^{k}\sum_{\ell_{2}\neq\ell_{1},j}^{k}\cdots\sum_{\ell_{\alpha}\neq\ell_{1},\ell_{2},\cdots,\ell_{\alpha-1},j}^{k}\sum_{\{k_{1},k_{2}\}\in B_{\ell_{1}}}\cdots\sum_{\{k_{\bar{\mathbf{q}}_{0}-1},k_{\bar{\mathbf{q}}_{0}}\}\in B_{\ell_{\alpha}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
    By definition of the dependnet structure, we have
@@
    $\leq$ $\left(\frac{p}{n}\right)^{\alpha}(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{\alpha}$
@@
    In this case, at least one cell consists of just one element of $C^{j}_{i1},\cdots,C^{j}_{i\bar{\mathbf{q}}_{0}}$, which implies we have $E\left[C^{j}_{ik_{1}}\cdots C^{j}_{i\bar{\mathbf{q}}_{0}}\right]=0$. Therefore, $A_{3}$ is equal to 0.

### v2

@@ D.3. Proof of Theorem 13
All the conditions for Theorem 3 and Theorem 10 trivially hold from Theorem 12 and assumptions of Theorem 12 as in Han and Shen (2023). ∎
D.4. Proof of Theorem 14
@@
All the conditions for Theorem 3 and Theorem 10 trivially hold from assumptions of Theorem 14 as in Theorem 13. ∎
Appendix E Proof of auxiliary results
E.1. Proof of Lemma 15
Proof.
Define $\bar{H}_{\beta}:=\beta\cdot\bar{H}$ and $F_{\beta}(A):=F_{\beta}(\bar{H}(w,A)_{w\in\mathcal{W}})$. As an initial step, we apply the result from Lemma 5, reducing the difference to the third derivatives of the functions involved. As shown in Appendix B.1, the third derivative of the function $g(n^{-1}F_{\beta}(\cdot))$ can be bounded by the derivatives of the loss function $\bar{H}$. For any $i\in\{1,\cdots,n\}$ and $j,k,\ell\in\{1,\cdots,d\}$, we have the following upper bound of derivatives of $\bar{H}$:
$|\langle\partial_{i,j}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}}\rangle_{\bar{H}_{\beta}}),$
$|\langle\partial_{i,j}\partial_{i,k}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{2}_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{2}}\rangle_{\bar{H}_{\beta}}),$
$|\langle\partial_{i,j}\partial_{i,k}\partial_{i,\ell}\bar{H}\rangle_{\bar{H}_{\beta}}|$ $\leq L^{3}_{n}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{3}}\rangle_{\bar{H}_{\beta}}),$
where we define $\bar{H}^{(i)}(w,A):=\ell_{i:\rho}((Aw)_{i})\geq 0$. so $\bar{H}(w,A)=\sum_{i=1}^{n}\bar{H}^{(i)}(w,A)+nR(w)$. Therefore, any $i\in\{1,\cdots,n\}$ and $j,k,\ell\in\{1,\cdots,d\}$, we have the following bounds:
@@ E.1. Proof of Lemma 15
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}}),$
@@
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}}),$
@@
$\leq 2^{2}L^{3}_{n}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{3q_{1}}\rangle_{\bar{H}_{\beta}}),$
@@
$\lesssim L_{n}^{3}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}+q_{2}}\rangle_{\bar{H}_{\beta}}),$
@@
$\lesssim L_{n}^{3}D^{3}_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{1}+q_{2}}\rangle_{\bar{H}_{\beta}}),$
@@
$\leq L_{n}^{3}D_{\ell}(\rho)(1+\langle\bar{H}^{(i)}(w,A)^{q_{3}}\rangle_{\bar{H}_{\beta}}.$
@@
$|\partial_{i,j}\partial_{i,k}\partial_{i,\ell}g(n^{-1}F_{\beta}(A))|\lesssim K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+\langle\bar{H}^{(i)}(w,A)^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}\right],$
where we define $\mathbf{q}=\max\{3q_{1},2q_{1}+2q_{2},q_{3}\}$. So, for any $j$ and $\ell,\ell^{\prime},\ell^{\prime\prime}\in B_{j}$, we obtain
$E[|\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))||X^{(B_{j})}_{i}$]
[…]
By using Chebyshev’s association inequality, we have
$E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta}}|X^{(B_{j})}_{i}]\leq E[\langle(\bar{H}^{(i)}(w,\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))^{\mathbf{q}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}$],
@@
[…]
@@
[…]
where we define $\mathbf{q}_{0}=q_{0}\mathbf{q}$ and $\bar{\mathbf{q}}_{0}:=2\lceil\mathbf{q}_{0}/2\rceil$, respectively. By trivial calculation, for any $1\leq j\leq k$ we have
$|t(X^{(B_{j})}_{i})^{\top}w|^{\mathbf{q}_{0}}\leq t^{\mathbf{q}_{0}}L^{\mathbf{q}_{0}}_{n}\left(\sum_{m\in B_{j}}|X_{i,m}|\right)^{\mathbf{q}_{0}}.$ (37)
@@
$\begin{split}E\left[\langle|((\mathbf{Z}_{i,j}^{0})^{\top}w)_{i}|^{\bar{\mathbf{q}}_{0}}\rangle_{\bar{H}_{\beta;-i}}|X^{(B_{j})}_{i}\right]\leq&E\left[\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\mathbf{q}_{0}}}\langle w_{k_{1}}\cdots w_{k_{\bar{\mathbf{q}}_{0}}}\rangle_{\bar{H}_{\beta;-i}}\right]\\ \leq&\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]E\left[\langle w_{k_{1}}\cdots w_{k_{\bar{\mathbf{q}}_{0}}}\rangle_{\bar{H}_{\beta;-i}}\right]\\ \leq&L_{n}^{\bar{\mathbf{q}}_{0}}\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right],\end{split}$ (38)
@@
$\begin{split}&E[|\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))||X^{(B_{j})}_{i}]\\ &\lesssim K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)\left[1+L^{\mathbf{q}}_{\ell_{i}}L_{n}^{\bar{\mathbf{q}}_{0}}\left(1+\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]+t^{\bar{\mathbf{q}}_{0}}\left(\sum_{m\in B_{j}}|X_{i,m}|\right)^{\bar{\mathbf{q}}_{0}}\right)\right].\end{split}$ (39)
The second term expresses the sum of all combinations of $\mathbf{{\bar{q}}}_{0}$ random variables exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$. Combining Lemma 5 with (39), we attain
$\left|E\left[\int^{1}_{0}(1-t)^{2}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime}\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}\partial_{i,\ell}\partial_{i,\ell^{\prime}}\partial_{i,\ell^{\prime\prime}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))dt\right]\right|$
$\leq E\left[\int^{1}_{0}(1-t)^{2}\left(\sum_{\ell\in B_{j}}\left|X_{i,\ell}\right|\right)^{3}E\left[\left|\partial_{i,(B_{j})_{\ell}}\partial_{i,(B_{j})_{\ell^{\prime}}}\partial_{i,(B_{j})_{\ell^{\prime\prime}}}g(n^{-1}F_{\beta}(\mathbf{Z}_{i,j}^{0}+tX^{(B_{j})}_{i}))\right|X^{(B_{j})}_{i}\right]dt\right]$
$\lesssim K_{g}n^{-1}\beta^{2}L_{n}^{3}D^{3}_{\ell}(\rho)E\left[\left(\sum_{\ell\in B_{j}}\left|X_{i,\ell}\right|\right)^{3}\right]$
$+K_{g}n^{-1}\beta^{2}L_{n}^{\bar{\mathbf{q}}_{0}+3}D^{3}_{\ell}(\rho)L^{\mathbf{q}}_{\ell_{i}}E\left[\left(\sum_{\ell\in B_{j}}\left|X_{i,\ell}\right|\right)^{3}\left(1+\left(\sum_{\ell\in B_{j}}\left|X_{i,\ell}\right|\right)^{\bar{\mathbf{q}}_{0}}+\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right)\right].$
@@ E.2. Proof of Lemma 16
[…]
@@
    4. Step 4.
      Set $\gamma_{2,\ell}$ as the number of variables included in the same dependent index cell as $C^{j}_{ik_{(\gamma_{1,\ell})+1}}$. As in Step 2, the possible ways to select variables are $d^{(\gamma_{2,\ell})-1}$. Denote those variables as $C^{j}_{ik_{(\gamma_{1,\ell})+2}},\cdots,C^{j}_{ik_{(\gamma_{1,\ell})+(\gamma_{2,\ell})}}$.
    5. Step 5.
@@
    $N_{1}$ $\lesssim_{q}\sum_{\ell=1}^{\bar{\mathbf{q}}_{0}-\alpha-1}p^{\ell}d^{\sum_{\ell^{*}=1}^{\ell}(\gamma_{\ell^{*},\ell}-1)}=\sum_{\ell=1}^{\bar{\mathbf{q}}_{0}-\alpha-1}p^{\ell}d^{\bar{\mathbf{q}}_{0}-\ell},$
    where the last equality follows from the definition of $\gamma_{\ell^{*},\ell}$, i.e. $\sum_{\ell^{*}=1}^{\ell}\gamma_{\ell^{*},\ell}=\bar{\mathbf{q}}_{0}$. Moreover, because $d\leq p$ holds, for any $1\leq\ell\leq\bar{\mathbf{q}}_{0}-\alpha-2$, we have
@@
    $=\sum_{\begin{subarray}{c}\ell_{1}=1\\ \ell_{1}\neq j\end{subarray}}^{k}\sum_{\ell_{2}\neq\ell_{1},j}^{k}\cdots\sum_{\ell_{\alpha}\neq\ell_{1},\ell_{2},\cdots,\ell_{\alpha-1},j}^{k}\sum_{\{k_{1},k_{2}\}\in B_{\ell_{1}}}\cdots\sum_{\{k_{\bar{\mathbf{q}}_{0}-1},k_{\bar{\mathbf{q}}_{0}}\}\in B_{\ell_{\alpha}}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right].$
    By definition of the dependent structure, we have
@@
    $\leq\left(\frac{p}{n}\right)^{\alpha}(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{\alpha},$
@@
    In this case, at least one cell consists of just one element of $C^{j}_{i1},\cdots,C^{j}_{i\bar{\mathbf{q}}_{0}}$, which implies we have $E[C^{j}_{ik_{1}}\cdots C^{j}_{i\bar{\mathbf{q}}_{0}}]=0$. Therefore, $A_{3}$ is equal to 0.